# Antisymmetric SU(2) adjoint generator as the universal origin of geometric phases in quantum two-level systems and classical polarization

### v2

@@ Abstract
We establish that the antisymmetric part of the adjoint SU(2) generator serves as the universal algebraic origin of geometric phases across classical polarization optics and quantum two-level systems. For classical ideal retarders, we demonstrate that the antisymmetric block of the Mueller matrix exclusively determines the geometric phase, encoding the angular-velocity pseudovector governing Stokes-vector evolution on the Poincaré sphere. Remarkably, the symmetric component is geometrically neutral. Quantum mechanically, the identical antisymmetric generator emerges in the adjoint SU(2) action on the Bloch sphere and fully governs the geometric phase of pure qubit states, independent of adiabaticity or cyclicity. This unified framework provides direct operational access to geometric phases via measured Mueller matrices or quantum process tomography, enabling precise control and diagnosis of geometric effects in both classical and quantum domains.
I Introduction
Geometric phases appear in a wide variety of wave phenomena, from classical polarization optics to quantum two-level dynamics (see, e.g., Ref. [1] for a recent review). In classical optics, Pancharatnam’s seminal work on interference between different polarization states introduced a geometric connection on the space of polarization states and led to what is now known as the Pancharatnam–Berry (PB) phase [2, 6]. In quantum mechanics, Berry showed that a quantum system undergoing adiabatic and cyclic evolution acquires a geometric phase that depends solely on the path traced in parameter space [3]. For two-level systems, these classical and quantum viewpoints share the same SU(2) and Bloch/Poincaré-sphere structure. This concept was subsequently generalized to nonadiabatic and noncyclic evolutions by Aharonov and Anandan [4] and by Samuel and Bhandari [5]. For polarized light, a comprehensive review of geometric and topological phase phenomena is given in Ref. [6]. Further generalizations to mixed quantum states in an interferometric setting were developed by Sjöqvist et al. [7].
@@ I Introduction
Here we make this correspondence explicit at the algebraic level and reveal a fundamental structure: the antisymmetric part of the adjoint $\mathrm{SU}(2)$ generator provides a universal kernel for geometric phases in both classical polarization optics and quantum two-level systems. This perspective not only unifies these seemingly disparate phenomena but also provides direct operational access to geometric phases from experimental measurements. On the classical side, we consider arbitrary ideal retarders described by Mueller matrices and show that their geometric action is completely encoded in the antisymmetric $3\times 3$ block acting on the Stokes vector. This antisymmetric block encodes the angular-velocity pseudovector that governs the tangential component of the Stokes-vector motion on the Poincaré sphere and therefore fully determines the associated geometric phase. The complementary symmetric block may reshape the mapping between Stokes vectors, but it is geometrically neutral in the sense that it does not contribute to the geometric phase.
On the quantum side, we show that the same antisymmetric structure appears in the adjoint action of any $\mathrm{SU}(2)$ unitary operator on a qubit. The Bloch vector of a pure qubit state evolves as $\mathbf{r}(t)=R(t)\,\mathbf{r}(0)$, where $R(t)\in\mathrm{SO}(3)$ is a continuous one-parameter family of rotations and its instantaneous generator $A(t)$, defined by $\dot{R}(t)=A(t)R(t)$, is an antisymmetric matrix. We demonstrate that the geometric phase of a pure qubit state is fully determined by this antisymmetric generator, independently of whether the evolution is adiabatic or nonadiabatic, cyclic or noncyclic. The role of the antisymmetric generator is again to encode the instantaneous angular velocity of the Bloch vector and, more specifically, the tangential component of this velocity on the Bloch sphere, which is the only part that contributes to the geometric phase.
@@
These results identify a single algebraic structure underlying PB phases in classical polarization and geometric phases in quantum two-level systems: the antisymmetric part of the adjoint $\mathrm{SU}(2)$ generator. This viewpoint is also operational. In classical experiments, the relevant antisymmetric block can be extracted directly from a measured Mueller matrix of an ideal retarder. In quantum experiments, the same antisymmetric generator can be reconstructed from process tomography through the adjoint map acting on the Bloch vector. In both cases this yields a simple criterion for isolating geometric-phase contributions and for diagnosing transformations where geometric phases are controlled or suppressed.
@@
The remainder of the paper is organized as follows. In Sec. II we analyze the Mueller matrix of an arbitrary ideal retarder and show that its antisymmetric block is the unique generator of geometric phase on the Poincaré sphere. In Sec. III we connect this antisymmetric generator with the tangential angular velocity of the Stokes vector and with the solid angle enclosed by its trajectory. Section IV extends the formalism to quantum two-level systems by studying the adjoint action of $\mathrm{SU}(2)$ on the Bloch sphere. In Sec. V we present illustrative classical and quantum examples that highlight the operational content of the theory. Section VI discusses possible extensions to more general quantum channels and higher-dimensional systems, and Sec. VII summarizes our conclusions.
II Classical polarization: Mueller retarders and antisymmetric generator
@@
We first consider the classical case of an ideal retarder acting on a fully polarized beam. The retarder is described in the Jones formalism by a unitary $2\times 2$ operator of the form [8]
@@
where $\delta$ is the retardance, $\mathbf{n}$ is a real unit vector that specifies the eigenpolarizations on the Poincaré sphere, and $\bm{\sigma}=(\sigma_{1},\sigma_{2},\sigma_{3})$ are the Pauli matrices. The pair $(\delta,\mathbf{n})$ provides a complete parametrization of pure retarders: $\delta$ determines the rotation angle on the Poincaré sphere and $\mathbf{n}$ the rotation axis.
The Pauli-algebraic approach to polarization optics was developed by Whitney in her early work on Pauli-algebraic operators in polarization optics [8]. A systematic development of a vectorial, pure operatorial Pauli-algebraic formalism for device and state operators has been given by Tudor [9, 10], providing an alternative but closely related framework to the standard Jones and Mueller calculi for deterministic (also called nondepolarizing) systems.
@@
where $\mathbf{s}=(S_{1},S_{2},S_{3})^{\mathsf{T}}$ denotes the normalized Stokes vector of a fully polarized state. The matrix $R$ is the adjoint $\mathrm{SO}(3)$ representation of the $\mathrm{SU}(2)$ operator $U$ and takes the standard Rodrigues form (see, e.g., Ref. [11, 12])
@@
Equations (7) and (8) show that the action of a pure retarder on the Poincaré sphere can be decomposed into two algebraically distinct components: a symmetric mapping that deforms the distribution of Stokes vectors along and orthogonal to the eigenaxis $\mathbf{n}$, and an antisymmetric generator that produces a rotation around $\mathbf{n}$. As we will show in Sec. III, only the antisymmetric generator $A$ contributes to the geometric phase: it encodes the instantaneous angular velocity of the Stokes vector and, in particular, its tangential component on the sphere. The symmetric part $S$ may modify the detailed mapping of Stokes vectors, but it does not contribute to the accumulated geometric phase.
For any ideal retarder the antisymmetric $3\times 3$ block of the Mueller matrix is therefore $A=\sin\delta\,[\mathbf{n}]_{\times},$ and this matrix will be identified in the following sections as the universal classical kernel of geometric phase.
@@
III Geometric phase and tangential angular velocity on the Poincaré sphere
The decomposition (6)–(8) shows that the polarimetric action of an ideal retarder on the Poincaré sphere is governed by an antisymmetric matrix $A$ of the form
$A=[\bm{\Omega}]_{\times},\qquad\bm{\Omega}=\sin\delta\,\mathbf{n},$ (9)
@@ III Geometric phase and tangential angular velocity on the Poincaré sphere
where $\Omega_{\mathrm{solid}}(C)$ is the oriented solid angle subtended by $C$ at the center of the sphere (see Fig. 1). The expression (12) can be derived from the Pancharatnam connection $\gamma=\arg\langle\psi(0)|\psi(T)\rangle-\text{(dynamical phase)}$ by recognizing that the overlap between initial and final states on the sphere is governed by the solid angle enclosed by the trajectory. For a comprehensive geometric treatment see [6, 5].
@@
The local geometry of these quantities is illustrated in Fig. 2. At a given time $t$, the unit Stokes/Bloch vector $\mathbf{s}(t)$ is normal to the sphere and therefore normal to the tangent plane $T_{\mathbf{s}}$ at that point. The instantaneous angular velocity vector $\bm{\Omega}(t)$ can be decomposed into a component parallel to $\mathbf{s}(t)$ and a component tangent to the sphere, $\bm{\Omega}(t)=\bm{\Omega}_{\parallel}(t)+\bm{\Omega}_{\perp}(t)$, with $\bm{\Omega}_{\parallel}(t)=(\bm{\Omega}\cdot\mathbf{s})\,\mathbf{s}$ and $\bm{\Omega}_{\perp}(t)=\bm{\Omega}(t)-\bm{\Omega}_{\parallel}(t)$. Here $\bm{\Omega}_{\parallel}$ is collinear with $\mathbf{s}$ and does not affect the motion on the sphere, whereas $\bm{\Omega}_{\perp}$ lies entirely in the tangent plane $T_{\mathbf{s}}$.
@@
Equation (11) shows that only $\bm{\Omega}_{\perp}$ contributes to the motion of $\mathbf{s}$, since
@@
This decomposition reveals why only the antisymmetric generator contributes to geometric phases: $\bm{\Omega}_{\parallel}$ is collinear with $\mathbf{s}$ and does not induce any motion on the sphere, whereas $\bm{\Omega}_{\perp}$ fully drives the evolution of $\mathbf{s}(t)$. In particular, the component of $\bm{\Omega}_{\perp}$ along the local normal to the osculating plane determines the geodesic curvature and thus the sign and magnitude of the geometric phase.
@@
Equations (9)–(19) establish that, for ideal retarders, the geometric phase is fully determined by the antisymmetric generator $A$. The symmetric part $S$ of the Mueller matrix does not enter in Eq. (11), and therefore it does not contribute to the solid angle or to $\gamma_{\mathrm{geom}}$. In other words, the antisymmetric $3\times 3$ block $A$ of the Mueller matrix contains the complete geometric content of the retarder: it specifies the instantaneous angular velocity $\bm{\Omega}$ and, in particular, its tangential component $\bm{\Omega}_{\perp}$ that controls the PB phase through Eq. (19).
@@
IV Quantum two-level systems: adjoint SU(2) action and Bloch-sphere geometric phase
@@
where $\mathbf{r}\in\mathbb{R}^{3}$ is the Bloch vector, $|\mathbf{r}|\leq 1$, and $\bm{\sigma}$ are the Pauli matrices. Pure states correspond to $|\mathbf{r}|=1$, so that $\mathbf{r}$ lies on the Bloch sphere.
@@
identical in form to Eq. (11). In this picture $\mathbf{r}(t)$ is a unit vector on the Bloch sphere and $\bm{\Omega}(t)$ plays the role of its instantaneous angular velocity on the sphere, now determined by the underlying $\mathrm{SU}(2)$ Hamiltonian, in direct analogy with the Stokes-vector evolution discussed in Sec. III.
@@
where $C$ is the closed Bloch trajectory and $\Omega_{\mathrm{solid}}(C)$ is the corresponding solid angle. This expression holds independently of whether the evolution is adiabatic or nonadiabatic and reduces to Berry’s original formula in the adiabatic case.
By the same geometric arguments as in Sec. III, the solid angle can be expressed in terms of the tangential component $\bm{\Omega}_{\perp}(t)$ of the angular velocity of $\mathbf{r}(t)$ on the sphere. In particular, one can write
@@
where $v(t)=|\dot{\mathbf{r}}(t)|$ is the speed of the Bloch vector along $C$ and $\kappa_{\mathrm{g}}(t)$ is the geodesic curvature of the trajectory. The geodesic curvature depends only on the tangential component $\bm{\Omega}_{\perp}(t)$ of the angular velocity and vanishes for geodesic motion on the sphere. Equation (32) therefore shows that the geometric phase of a pure qubit state is entirely determined by the antisymmetric generator $A_{U}(t)$.
From the operational point of view, this means that the geometric content of a unitary qubit evolution can be extracted solely from the antisymmetric part of the adjoint map $R_{U}(t)$. In experiments where the quantum process is characterized via qubit process tomography, one reconstructs the completely positive trace-preserving map $\mathcal{E}$ on qubit density matrices $\rho$, i.e., the map $\rho\mapsto\mathcal{E}(\rho)$, and the induced action on the Bloch vector,

### v3

# Antisymmetric Mueller generator as the universal origin of geometric phase in classical polarization and quantum two-level systems

@@ Abstract
We show that the antisymmetric Mueller generator provides a universal algebraic kernel for geometric phase in classical polarization optics and in quantum two-level systems. For any ideal retarder, the antisymmetric $3\times 3$ block of its Mueller matrix—the antisymmetric generator of the adjoint $\mathrm{SU}(2)$ action on the Stokes vector—encodes the angular-velocity vector that drives the tangential motion on the Poincaré sphere and fully determines the Pancharatnam–Berry phase, while the symmetric block is geometrically neutral. The same antisymmetric generator governs the evolution of pure qubit states on the Bloch sphere. This unified viewpoint yields operational criteria to identify and control geometric-phase contributions from measured Mueller matrices and from qubit process tomography.
1 Introduction
Geometric phase is a ubiquitous concept in modern physics, with manifestations ranging from optics to quantum mechanics and condensed matter [1, 2]. In classical polarization optics, the Pancharatnam–Berry (PB) phase arises whenever the polarization state traces a closed contour on the Poincaré sphere [3, 4]. This effect underlies a wide variety of polarization-based geometric-phase elements, from geometric-phase lenses to wavefront-shaping metasurfaces, whose operation is routinely described and characterized within the Mueller-matrix framework and related polarimetric formalisms [5, 6, 7].
Several algebraic formulations have been developed to evaluate the PB phase introduced by general polarization devices described by Jones matrices, including systems with nonorthogonal eigenpolarizations [8, 9]. These approaches focus on the overall geometric phase associated with a given optical system, typically in terms of its Jones representation and the associated paths on the Poincaré sphere.
In quantum mechanics, Berry showed that a system undergoing adiabatic and cyclic evolution acquires a geometric phase that depends only on the path in parameter space [2]. This concept was subsequently generalized to nonadiabatic and noncyclic evolutions by Aharonov and Anandan [10] and by Samuel and Bhandari [11]. For a two-level system, pure states can be represented by points on the Bloch sphere, and the associated geometric phases are again determined by solid angles on that sphere. For polarized light, a comprehensive review of geometric and topological phase phenomena is given in Ref. [4]. Further generalizations to mixed quantum states in an interferometric setting were developed by Sjöqvist et al. [12].
@@ 1 Introduction
Beyond reproducing standard formulas for geometric phases in these familiar settings, the aim of this work is to present an explicitly algebraic and operational reformulation in which a single antisymmetric adjoint generator is identified as the common carrier of geometric-phase information in both regimes, and can be accessed directly from experimentally measured quantities
A practical point is that a Mueller (or adjoint) matrix describes an one-step state transformation, whereas a geometric phase is associated with a path on the Poincaré/Bloch sphere. The connection is local: the antisymmetric block yields the angular-velocity generator that drives the tangential motion of the Stokes/Bloch vector. When a device parameter varies continuously (e.g., along propagation in a birefringent medium or under rotation of a wave plate), or when a sequence of such transformations is applied, the state traces a curve; the accumulated geometric phase then follows from the solid angle enclosed by that curve. In this sense, identifying $A$ is a differential statement: it fixes the instantaneous tangential velocity and thereby the geometric phase accumulated along any path.
@@
In this work we make this correspondence algebraically explicit and operational. We identify the antisymmetric Mueller generator, namely the antisymmetric part of the adjoint $\mathrm{SU}(2)$ action on the Stokes vector, as the universal kernel for geometric phase in both classical polarization optics and quantum two-level systems. On the classical side, we show that for any ideal retarder the antisymmetric $3\times 3$ block of its Mueller matrix completely encodes the angular-velocity pseudovector that drives the tangential component of the Stokes-vector motion and therefore fully determines the Pancharatnam–Berry phase, while the complementary symmetric block is geometrically neutral. On the quantum side, we show that the same antisymmetric generator arises in the adjoint action of $\mathrm{SU}(2)$ unitaries on the Bloch sphere and fully determines the geometric phase of pure qubit states, independently of adiabaticity or cyclicity. The role of the antisymmetric generator is again to encode the instantaneous angular velocity of the Bloch vector and, more specifically, the tangential component of this velocity on the Bloch sphere, which is the only part that contributes to the geometric phase.
@@
These results identify a single algebraic structure underlying PB phases in classical polarization and geometric phases in quantum two-level systems: the antisymmetric part of the adjoint $\mathrm{SU}(2)$ generator. Beyond providing a unified conceptual framework, this viewpoint is operational. In classical experiments, the antisymmetric Mueller generator can be extracted directly from a measured Mueller matrix of an ideal retarder. In quantum experiments, the same antisymmetric generator can be reconstructed from process tomography through the adjoint map acting on the Bloch vector. In both cases this yields a practical criterion for isolating geometric-phase contributions and for designing or diagnosing transformations where geometric phases are controlled or suppressed.
@@
The remainder of the paper is organized as follows. In Sec. 2 we analyze the Mueller matrix of an arbitrary ideal retarder and show that its antisymmetric block is the unique generator of geometric phase on the Poincaré sphere. In Sec. 3 we connect this antisymmetric generator with the tangential angular velocity of the Stokes vector and with the solid angle enclosed by its trajectory. Section 4 extends the formalism to quantum two-level systems by studying the adjoint action of $\mathrm{SU}(2)$ on the Bloch sphere. In Sec. 5 we present illustrative classical and quantum examples that highlight the operational content of the theory. Section 6 discusses possible extensions to more general quantum channels and higher-dimensional systems, and Sec. 7 summarizes our conclusions.
2 Classical polarization: Mueller retarders and antisymmetric generator
@@
We first consider the classical case of an ideal retarder acting on a fully polarized beam. Here “ideal retarder” means a lossless, nondepolarizing, nondiattenuating element that introduces a relative phase delay between two orthogonal eigenpolarizations, without changing total intensity or degree of polarization. The retarder is described in the Jones formalism by a unitary $2\times 2$ operator of the form [13]
@@
where $\delta$ is the retardance, $\mathbf{n}$ is a real unit vector that specifies the eigenpolarizations on the Poincaré sphere, and $\bm{\sigma}=(\sigma_{1},\sigma_{2},\sigma_{3})$ is the Pauli vector, i.e., the vector of Pauli matrices. The pair $(\delta,\mathbf{n})$ provides a complete parametrization of pure retarders: $\delta$ determines the rotation angle on the Poincaré sphere and $\mathbf{n}$ the rotation axis.
The Pauli-algebraic approach to polarization was developed by Whitney in her early work on Pauli operators [13]. A systematic development of a vectorial, pure operatorial Pauli-algebraic formalism for device and state operators has been given by Tudor [14, 15], providing an alternative but closely related framework to the standard Jones and Mueller calculi for deterministic (also called nondepolarizing) systems.
@@
where $\mathbf{s}=(S_{1},S_{2},S_{3})^{\mathsf{T}}$ denotes the normalized Stokes vector of a fully polarized state. The matrix $R$ is the adjoint $\mathrm{SO}(3)$ representation of the $\mathrm{SU}(2)$ operator $U$ and takes the standard Rodrigues form (see, e.g., Refs. [16, 17])
@@
At the infinitesimal level relevant to geometric phase, only the antisymmetric part acts as a generator of motion on the sphere (via a cross product with an angular-velocity vector), whereas the symmetric part does not generate tangential transport and therefore cannot contribute to the swept solid angle.
@@
Equations (7) and (8) show that the action of a pure retarder on the Poincaré sphere splits into two algebraically distinct components: a symmetric mapping that reshapes the distribution of Stokes vectors relative to the eigenaxis $\mathbf{n}$, and a rotational component governed by the antisymmetric block $A$. As will be discussed in Sec. 3, it is this rotational component that contains the geometric-phase information, whereas the symmetric part affects the mapping of Stokes vectors without contributing to the accumulated phase.
For any ideal retarder the antisymmetric $3\times 3$ block of the Mueller matrix is therefore
$A=\sin\delta\,[\mathbf{n}]_{\times},$ (9)
and this matrix will be identified in the following sections as the universal classical kernel of geometric phase.
@@
3 Geometric phase and tangential angular velocity on the Poincaré sphere
The decomposition (6)–(8) shows that, although the complete polarimetric action of an ideal retarder on the Poincaré sphere is described by the rotation $R$, the geometric-phase behavior is governed by an antisymmetric matrix $A=[\bm{\Omega}]_{\times}$ with
$\bm{\Omega}=\sin\delta\,\mathbf{n},$ (10)
@@ 3 Geometric phase and tangential angular velocity on the Poincaré sphere
where $\Omega_{\mathrm{solid}}(C)$ is the oriented solid angle subtended by $C$ at the center of the sphere (see Fig. 1). The expression (13) can be derived from the Pancharatnam connection
$\gamma=\arg\langle\psi(0)|\psi(T)\rangle-\gamma_{\rm dyn},$ (14)
where $\arg\langle\psi(0)|\psi(T)\rangle$ is the total phase difference between the initial and final states and $\gamma_{\rm dyn}$ is the dynamical phase. The overlap between initial and final states on the sphere is governed by the solid angle enclosed by the trajectory. For a comprehensive geometric treatment see [4, 11].
@@
Equivalently, one may view Eq. (13) as a direct consequence of the Gauss–Bonnet theorem on the sphere [23]: for a closed loop completed by geodesic segments (as in Pancharatnam’s geodesic closure), the enclosed oriented area is tied to the integrated geodesic curvature and fixes the geometric phase.
@@
The local geometry of these quantities is illustrated in Fig. 2. At a given time $t$, the unit Stokes/Bloch vector $\mathbf{s}(t)$ is normal to the sphere and therefore normal to the tangent plane $T_{\mathbf{s}}$ at that point.
@@
From a more intuitive viewpoint, the component of the angular velocity that is parallel to the Stokes/Bloch vector merely changes an overall phase convention and leaves the point fixed on the sphere, so it cannot contribute to the swept solid angle, whereas only the component that lies in the tangent plane actually displaces the state along a curve and is therefore responsible for the geometric phase.
Equation (12) then shows that the motion of $\mathbf{s}(t)$ on the sphere is driven entirely by $\bm{\Omega}_{\perp}$, since
@@
Thus $\bm{\Omega}_{\parallel}$ is collinear with $\mathbf{s}$ and does not induce any motion on the sphere, whereas $\bm{\Omega}_{\perp}$ fully drives the evolution of $\mathbf{s}(t)$. In particular, the component of $\bm{\Omega}_{\perp}$ along the local normal to the osculating plane determines the geodesic curvature of the trajectory and hence the sign and magnitude of the geometric phase.
@@
Taken together, Eqs. (10)–(21) show that, for ideal retarders, the evolution relevant to geometric phase is governed entirely by the antisymmetric part of the Mueller block. The symmetric contribution does not appear in the kinematic description of the motion on the sphere and therefore plays no role in determining the solid angle or the geometric phase. In other words, the antisymmetric $3\times 3$ block $A$ of the Mueller matrix contains the complete geometric-phase content of the retarder: it specifies the instantaneous angular velocity $\bm{\Omega}$ and, in particular, its tangential component $\bm{\Omega}_{\perp}$ that controls the PB phase through Eq. (21).
@@
4 Quantum two-level systems: adjoint SU(2) action and Bloch-sphere geometric phase
@@
Working in the Bloch-vector representation is automatically gauge invariant, because the dynamics is encoded in density matrices and real three-dimensional trajectories, so no explicit choice of phase for the underlying state vector is required; the geometric phase is directly associated with the geometry of the path on the Bloch sphere.
@@
where $\mathbf{r}\in\mathbb{R}^{3}$ is the Bloch vector, $|\mathbf{r}|\leq 1$, and $\bm{\sigma}$ is the Pauli vector. Pure states correspond to $|\mathbf{r}|=1$, so that $\mathbf{r}$ lies on the Bloch sphere.
@@
identical in form to Eq. (12). In this picture $\mathbf{r}(t)$ is a unit vector on the Bloch sphere and $\bm{\Omega}(t)$ plays the role of its instantaneous angular velocity on the sphere, now determined by the underlying $\mathrm{SU}(2)$ Hamiltonian, in direct analogy with the Stokes-vector evolution discussed in Sec. 3.
@@
where $C$ is the closed Bloch trajectory and $\Omega_{\mathrm{solid}}(C)$ is the corresponding solid angle. This expression holds independently of whether the evolution is adiabatic or nonadiabatic and reduces to Berry’s original formula in the adiabatic case. It is precisely the Aharonov–Anandan geometric phase for cyclic evolutions of pure states and therefore provides a nonadiabatic extension of Berry’s result within the same geometric framework.
By the same geometric arguments as in Sec. 3, the solid angle can be expressed in terms of the tangential component $\bm{\Omega}_{\perp}(t)$ of the angular velocity of $\mathbf{r}(t)$ on the sphere. In particular, one can write
@@
where $v(t)=|\dot{\mathbf{r}}(t)|$ is the speed of the Bloch vector along $C$ and $\kappa_{\mathrm{g}}(t)$ is the geodesic curvature of the trajectory. The geodesic curvature depends only on the tangential component $\bm{\Omega}_{\perp}(t)$ of the angular velocity and vanishes for geodesic motion on the sphere. Equation (34) shows that, for pure-state qubit evolutions, the geometric contribution to the phase is encoded in the antisymmetric part of the adjoint map through the tangential angular velocity of the Bloch vector. The remaining components of the transformation do not influence the enclosed solid angle and hence do not affect the geometric phase.
From the operational point of view, this means that the geometric-phase content of a unitary qubit evolution can be extracted solely from the antisymmetric part of the adjoint map $R_{U}(t)$. In experiments where the quantum process is characterized via qubit process tomography, one reconstructs the completely positive trace-preserving map $\mathcal{E}$ on qubit density matrices $\rho$, i.e., the map $\rho\mapsto\mathcal{E}(\rho)$, and the induced action on the Bloch vector,